On 1/1/22 Big Co acquired 60% of Little Co voting stock for $300,000; the fair value of the non-controlling interest was $200,000 on that date.

Little's book value on that date was $350,000. Little had the following misvalued/unreported assets and liabilities:

Land: Undervalued by $5,000

Inventory, FIFO basis: Undervalued by $12,000

Bonds payable, 5-year life: Undervalued by $10,000

In-process R&D, not reported on Little's balance sheet, 2-year life: Worth $8,000

In 2022 Little reported earnings of $50,000 and paid dividends of $10,000

What is Little's "adjusted income" after taking into account the various differential amortizations?
Answer #1
Book value 350000
Land 5000
inventory 12000
Bonds payable -10000
In process R&D 8000
365000
Purchase price 300000
NCI 200000
Goodwill 135000
Goodwill Big Co 81000
Goodwill NCI 54000
Amortisation/interest differentail
Bonds =excess of fairvalue over book value/useful life
=-10000/5
-2000
inprocess R&D =excess of fairvalue over book value/useful lofe
=8000/2
4000
Differential mortisation =4000-2000
2000
Adjusted income
=net income-differential amortisation
=50000-2000
48000
Share in income
Big Co 28800
NCI 19200
Investment Value
Initial investment 300000
Add: adjusted income 28800
Less: share of dividend -6000
322800